# A Family of Berndt-Type Integrals and Associated Barnes Multiple Zeta Functions

### v1

@@ 5 Evaluations of Hyperbolic Summations via Jacobi Functions
$Y_{2n+1,1}\left(y\right)=\displaystyle\sum_{j=1}^{\infty}{\displaystyle\frac{(-1)^{j}\left(2j-1\right)^{2n}}{\sinh\left(\displaystyle\frac{2j-1}{2}y\right)}}=(-1)^{n}z^{2n+1}\sqrt{x}\displaystyle\frac{\mathrm{S}_{2n+1}\left(x\right)}{2}.$ (5.55)
Differentiating (5.55) gives the relation
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}Y_{p,1}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{(-1)^{n}\left(2n-1\right)^{p-1}}{\sinh\left(\displaystyle\frac{2n-1}{2}y\right)}=-\displaystyle\frac{1}{2}}\text{DY}_{p,2}\left(y\right).$
Using the transformation $\frac{dx}{dy}=-x(1-x)z^{2}$ ([5, P. 120, Entry. 9(i)] or (2.11)), we obtain
$\text{DY}_{p,2}\left(y\right)=2x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}Y_{p,1}\left(y\right).$ (5.56)
Applying Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we have
@@
$Y_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)=-(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p}\displaystyle\frac{\mathrm{S}_{p-1}\left(1-x\right)}{2}\sqrt{1-x},$ (5.57)
@@
$\text{DY}_{p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}Y_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}Y_{p,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.58)
What’s more, as known from Rui-Xu-Zhao’s paper[17, Thm. 3.15] and apply Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we have
@@
Finally, combining equations (5), (5.55), (5.57), (5) and (5.59), we thus complete the proof of Theorem 5.1. ∎
@@
By Mathematica, we have
@@
.$\Big{.}\qquad\qquad\qquad+256(x-1)\left(1382x^{4}-2764x^{3}+1641x^{2}-259x+2\right)\Big{.}$
.$\Big{.}\quad-\sqrt{1-x}\left(555731x^{5}-1528258x^{4}+1467472x^{3}-562528x^{2}+68096x-512\right)z\Big{\}},$
@@
.$\Big{.}\quad\qquad\qquad\qquad+2048(x-1)\left(929569x^{6}-2788707x^{5}+3021099x^{4}\atop-1394353x^{3}+240594x^{2}-8202x+4\right)\Big{.}$
.$\Big{.}\quad-\sqrt{1-x}\left(2990414715x^{7}-11214055178x^{6}+16307105232x^{5}-11436042048x^{4}\atop+3872630528x^{3}-536879616x^{2}+16834560x-8192\right)z\Big{\}}.$
@@
Set $\Gamma=\Gamma(1/4)$ and $p=4m-1$. Then, for any integer $m>0$, we know that
@@
Set $\Gamma=\Gamma(1/4)$. By Mathematica, we have
@@
$G_{p,2}^{\prime}\left(y\right)$ $=-\displaystyle\frac{1}{2}z^{p}\mathrm{A}_{p-1}(1-x)\sqrt{x}+(p-1)z^{p}z^{\prime}x(1-x)\mathrm{P}_{p-2}(1-x)\sqrt{x(1-x)}$
@@
where $\mathrm{A}_{p-1}\left(1-x\right)$ are coefficients in the Maclaurin expansion of $\mathrm{nd}\left(u\right)$ and $\mathrm{P}_{p-2}\left(1-x\right)$
are coefficients in the Maclaurin expansion of $\mathrm{sd}\left(u\right)$.
@@
Starting from equation (4.42), we derive
$G_{p,2}^{\prime}\left(y\right)$ $=-(-1)^{(p-1)/2}\displaystyle\frac{2^{p-1}\pi^{p}}{y^{p}}B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)+(-1)^{(p-1)/2}\displaystyle\frac{2(p-1)\pi^{p-1}}{y^{p}}X_{p-2,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$\quad-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+1}}{y^{p+1}}\text{DX}_{p-2,2}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.62)
Now, we need to compute the series $B_{p,1}\left(y\right)$. From Lemma 3.5 and [8], we obtain the Maclaurin expansion for $nd(u)$
$\mathrm{nd}(u)=\displaystyle\sum_{n=0}^{\infty}{\mathrm{A}_{2n}\left(x\right)\displaystyle\frac{\left(-1\right)^{n}u^{2n}}{\left(2n\right)!}},\quad\displaystyle\frac{2q^{n}}{1+q^{2n}}=\displaystyle\frac{1}{\cosh(ny)}.$ (5.63)
@@
Setting $q\equiv q\left(x\right)=e^{-y}$, we expand
${\rm nd}(u)$ $=\displaystyle\frac{\pi}{2Kk^{\prime}}+\displaystyle\frac{2\pi}{\mathrm{K}k^{\prime}}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{(-1)^{n}q^{n}n^{2j}}{1+q^{2n}}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{\mathrm{K}}\right)^{2j}}{\left(2j\right)!}}$
$=\displaystyle\frac{\pi}{2\mathrm{K}k^{\prime}}+\displaystyle\frac{\pi}{\mathrm{K}k^{\prime}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{\mathrm{K}}\right)^{2j}}{\left(2j\right)!}}B_{2j+1,1}\left(y\right).$ (5.65)
Comparing the coefficients of $u^{2n}$ in (5.63) and (5), we deduce
$B_{p,1}\left(y\right)=(-1)^{(p-1)/2}z^{p}\sqrt{1-x}\displaystyle\frac{\mathrm{A}_{p-1}\left(x\right)}{2^{p}}.$ (5.66)
Applying Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we obtain
$B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)=(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p}\sqrt{x}\displaystyle\frac{\mathrm{A}_{p-1}\left(1-x\right)}{2^{p}}.$ (5.67)
By using the results in Rui-Xu-Zhao’s paper[17, Thm. 3.3], we have
$X_{p,1}^{\prime}(y)=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\cosh\left(\displaystyle\frac{(2n-1)y}{2}\right)}=-(-1)^{(p-1)/2}z^{p+1}\sqrt{x(1-x)}\displaystyle\frac{\mathrm{P}_{p}(x)}{2}.$ (5.68)
Differentiating $X_{p-2,1}^{\prime}\left(y\right)$ yields
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}X_{p-2,1}^{\prime}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{(-1)^{n}\left(2n-1\right)^{p-2}}{\cosh\left(\displaystyle\frac{2n-1}{2}y\right)}=-\displaystyle\frac{1}{2}}\text{DX}_{p-2,2}^{\prime}\left(y\right).$
Using $\frac{dx}{dy}=-x(1-x)z^{2}$ ([5, P. 120, Entry. 9(i)] or (2.11)), we get
$\text{DX}_{p-2,2}^{\prime}\left(y\right)=2x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}X_{p-2,1}^{\prime}\left(y\right).$ (5.69)
Applying Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we obtain
@@
Hence, the (5.68) and (5.69) can be rewritten as
$X_{p-2,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)=(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p+1}\displaystyle\frac{\mathrm{P}_{p-2}\left(1-x\right)}{2}\sqrt{x(1-x)}$ (5.70)
@@
$\text{DX}_{p-2,2}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}X_{p-2,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}X_{p-2,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.71)
@@
By Mathematica, we have
@@
.$\Big{.}\quad+\left(-79360x^{5}+198400x^{4}-166112x^{3}+50768x^{2}-3698x+1\right)\sqrt{1-x}z\Big{.}$
.$\Big{.}\quad+50521x^{5}-138933x^{4}+130250x^{3}-45530x^{2}+3693x-1\Big{\}},$
@@
.$\Big{.}\quad+\left(-313155584x^{7}+1096044544x^{6}-1458129408x^{5}\atop+905212160x^{4}-255034240x^{3}+25361472x^{2}-298946x+1\right)\sqrt{(1-x)}z\Big{.}$
.$\Big{.}\quad+(199360981x^{7}-747603679x^{6}+1074680289x^{5}-728130163x^{4}\atop+226132303x^{3}-24738669x^{2}+298939x-1\Big{\}}.$
@@
Set $\Gamma=\Gamma(1/4)$ and $p=4m-1$. Then, for any integer $m>0$ we know that
@@
$\quad-\displaystyle\frac{1}{2^{4m}}(4m-2)\mathrm{P}_{4m-3}\left(\displaystyle\frac{1}{2}\right)\displaystyle\frac{1}{\sqrt{2}}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{6m-2}}.$ (5.72)
@@
Set $\Gamma=\Gamma(1/4)$. By Mathematica, we have
@@
Let $p\geq 3$ be an odd integer. We have
$\bar{G}_{p,1}\left(y\right)$ $=\displaystyle\frac{1}{2}z^{p+1}\mathrm{P}_{p-1}(1-x)\sqrt{x(1-x)}+pz^{p+1}z^{\prime}x(1-x)\mathrm{A}_{p-1}(1-x)\sqrt{x}$ (5.73)
@@
where $\mathrm{P}_{p}\left(1-x\right)$ denotes coefficients in the Maclaurin expansion of $\mathrm{sd}\left(u\right)$ and $\mathrm{A}_{p-1}\left(1-x\right)$
denotes coefficients in the Maclaurin expansion of $\mathrm{nd}\left(u\right)$.
@@
We begin with equation (4.43), which gives
$\bar{G}_{p,1}\left(y\right)$ $=-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+1}}{y^{p+1}}X_{p,1}^{\prime}\left(\displaystyle\frac{\pi^{2}}{y}\right)-(-1)^{(p-1)/2}\displaystyle\frac{p2^{p}\pi^{p}}{y^{p}}B_{p,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$\quad+(-1)^{(p-1)/2}\displaystyle\frac{2^{p}\pi^{p+2}}{y^{p+2}}\text{DB}_{p,2}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.74)
The derivative of $B_{p,1}^{\prime}\left(y\right)$ satisfies
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}B_{p,1}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{(-1)^{n}n^{p-1}}{\cosh(ny)}=-\text{DB}_{p,2}\left(y\right).$
From the fundamental relation $\frac{dx}{dy}=-x(1-x)z^{2}$ ([5, P. 120, Entry. 9(i)] or [17]), we have
$\text{DB}_{p,2}\left(y\right)=x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}B_{p,1}\left(y\right).$ (5.75)
Using Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we obtain

### v3

@@ 5 Evaluations of Hyperbolic Summations via Jacobi Functions
$\tilde{G}_{2n,1,0}\left(y\right)=\displaystyle\sum_{j=1}^{\infty}{\displaystyle\frac{(-1)^{j}\left(2j-1\right)^{2n}}{\sinh\left(\displaystyle\frac{2j-1}{2}y\right)}}=(-1)^{n}z^{2n+1}\sqrt{x}\displaystyle\frac{\mathrm{S}_{2n+1}\left(x\right)}{2}.$ (5.56)
Fix $\delta>0$. For $y\geq\delta$ we have
$\left|\displaystyle\frac{(-1)^{j}(2j-1)^{2n}}{\sinh\!\left(\displaystyle\frac{2j-1}{2}y\right)}\right|\leq\displaystyle\frac{(2j-1)^{2n}}{\sinh\!\left(\displaystyle\frac{2j-1}{2}\delta\right)}\leq 2(2j-1)^{2n}e^{-\tfrac{2j-1}{2}\delta},$
hence $\widetilde{G}_{2n,1,0}(y)$ converges uniformly on $\delta,\infty)$ by the Weierstrass M-test. Moreover,
$f_{j}(y):=\displaystyle\frac{(-1)^{j}(2j-1)^{2n}}{\sinh\!\left(\displaystyle\frac{2j-1}{2}y\right)}\quad\Rightarrow\quad f_{j}^{\prime}(y)=-\displaystyle\frac{(-1)^{j}(2j-1)^{2n+1}}{2}\,\displaystyle\frac{\cosh\!\left(\displaystyle\frac{2j-1}{2}y\right)}{\sinh^{2}\!\left(\displaystyle\frac{2j-1}{2}y\right)}.$
Using $\frac{\cosh t}{\sinh^{2}t}\leq Ce^{-t}$ for $t=\displaystyle\frac{2j-1}{2}y\geq\displaystyle\frac{2j-1}{2}\delta$, we obtain
$|f_{j}^{\prime}(y)|\leq C(2j-1)^{2n+1}e^{-\tfrac{2j-1}{2}\delta},$
so the derivative series $\sum_{j\geq 1}f_{j}^{\prime}(y)$ converges uniformly on $\delta,\infty)$. Hence, differentiating (5.56) gives the relation
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\tilde{G}_{p-1,1,0}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{(-1)^{n}\left(2n-1\right)^{p-1}}{\sinh\left(\displaystyle\frac{2n-1}{2}y\right)}=-\displaystyle\frac{1}{2}}\tilde{G}_{p,2,-1}\left(y\right).$
By applying the transformation $\frac{\mathrm{d}x}{\mathrm{d}y}=-x(1-x)z^{2}$ ([5, p. 120, Entry. 9(i)] or (2.11)), we obtain
$\tilde{G}_{p,2,-1}\left(y\right)=2x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\tilde{G}_{p-1,1,0}\left(y\right).$ (5.57)
Resorting to Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we have
@@
$\tilde{G}_{p-1,1,0}\left(\displaystyle\frac{\pi^{2}}{y}\right)=-(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p}\displaystyle\frac{\mathrm{S}_{p-1}\left(1-x\right)}{2}\sqrt{1-x},$ (5.58)
@@
$\tilde{G}_{p,2,-1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}\tilde{G}_{p-1,1,0}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\tilde{G}_{p-1,1,0}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.59)
What’s more, as known from Rui-Xu-Zhao’s paper[19, Thm. 3.14] and apply Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we deduce
@@
Finally, combining (5), (5.56), (5.58), (5) and (5.60), we thus complete the proof of Theorem 5.1. ∎
@@
Using Mathematica to evaluate specific instances of the right-hand side of (5.1), we derive the following examples:
@@
$\Big.\qquad\qquad\qquad+256(x-1)\left(1382x^{4}-2764x^{3}+1641x^{2}-259x+2\right)\Big$.
$\Big.\quad-\sqrt{1-x}\left(555731x^{5}-1528258x^{4}+1467472x^{3}-562528x^{2}+68096x-512\right)z\Big\},$
@@
$\Big.\quad\qquad\qquad\qquad+2048(x-1)\left(929569x^{6}-2788707x^{5}+3021099x^{4}\atop-1394353x^{3}+240594x^{2}-8202x+4\right)\Big$.
$\Big.\quad-\sqrt{1-x}\left(2990414715x^{7}-11214055178x^{6}+16307105232x^{5}-11436042048x^{4}\atop+3872630528x^{3}-536879616x^{2}+16834560x-8192\right)z\Big\}.$
Setting p=4m1p=4m-1 and $x=1/2$ in Theorem 5.1 yields the following corollary.
@@
Set $\Gamma=\Gamma(1/4)$ and $\mathrm{S}^{\prime}_{p}(x)=\mathrm{d}\mathrm{S}_{p}(x)/\mathrm{d}x$. For any integer $m>0$, we obtain
@@
Set $\Gamma=\Gamma(1/4)$. By substituting $y=\pi$ and $x=1/2$ into Example 5.2 and evaluating the result with Mathematica, we obtain the following examples:
@@
$\tilde{G}_{p-1,1,2}\left(y\right)$ $=-\displaystyle\frac{1}{2}z^{p}\mathrm{A}_{p-1}(1-x)\sqrt{x}+(p-1)z^{p}z^{\prime}x(1-x)\mathrm{P}_{p-2}(1-x)\sqrt{x(1-x)}$
@@
where $\mathrm{A}_{p-1}\left(x\right)$ are coefficients in the Maclaurin expansion of $\mathrm{nd}\left(u\right)$ and $\mathrm{P}_{p}\left(x\right)$ are coefficients in the Maclaurin expansion of $\mathrm{sd}\left(u\right)$ (see [19, Lem. 3.1]).
@@
To establish the result, we start with (4.43)
$\tilde{G}_{p-1,1,2}\left(y\right)$ $=-(-1)^{(p-1)/2}\displaystyle\frac{2^{p-1}\pi^{p}}{y^{p}}G_{p-1,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)+(-1)^{(p-1)/2}\displaystyle\frac{2(p-1)\pi^{p-1}}{y^{p}}\tilde{G}_{p-2,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$\quad-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+1}}{y^{p+1}}\tilde{G}_{p-1,-1,2}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.63)
At this point, we shall compute the series $G_{p-1,0,1}\left(y\right)$. From Lemma 3.5 and [10], we acquire the Maclaurin expansion for $\mathrm{nd}(u)$
$\mathrm{nd}(u)=\displaystyle\sum_{n=0}^{\infty}{\mathrm{A}_{2n}\left(x\right)\displaystyle\frac{\left(-1\right)^{n}u^{2n}}{\left(2n\right)!}}\quad\text{and}\quad\displaystyle\frac{2q^{n}}{1+q^{2n}}=\displaystyle\frac{1}{\cosh(ny)},$ (5.64)
@@
Setting $q\equiv q\left(x\right):=e^{-y}$, we expand
${\rm nd}(u)$ $=\displaystyle\frac{\pi}{2Kk^{\prime}}+\displaystyle\frac{2\pi}{{K}k^{\prime}}\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{(-1)^{n}q^{n}n^{2j}}{1+q^{2n}}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{{K}}\right)^{2j}}{\left(2j\right)!}}$
$=\displaystyle\frac{\pi}{2{K}k^{\prime}}+\displaystyle\frac{\pi}{{K}k^{\prime}}\displaystyle\sum_{j=0}^{\infty}{\displaystyle\frac{\left(-1\right)^{j}\left(\displaystyle\frac{\pi u}{{K}}\right)^{2j}}{\left(2j\right)!}}G_{2j,0,1}\left(y\right).$ (5.66)
Through a comparison of the coefficients of $u^{2n}$ in (5.64) and (5), we derive
$G_{p-1,0,1}\left(y\right)=(-1)^{(p-1)/2}z^{p}\sqrt{1-x}\displaystyle\frac{\mathrm{A}_{p-1}\left(x\right)}{2^{p}}.$ (5.67)
Employing Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we obtain
$G_{p-1,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)=(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p}\sqrt{x}\displaystyle\frac{\mathrm{A}_{p-1}\left(1-x\right)}{2^{p}}.$ (5.68)
By using the results in Rui-Xu-Zhao’s paper[19, Thm. 3.2], we have
$\tilde{G}_{p,0,1}(y)=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\cosh\left(\displaystyle\frac{(2n-1)y}{2}\right)}=(-1)^{(p-1)/2-1}z^{p+1}\sqrt{x(1-x)}\displaystyle\frac{\mathrm{P}_{p}(x)}{2}.$ (5.69)
Consider the series $\tilde{G}_{p,0,1}(y)$, and fix $\delta>0$. For $y\geq\delta$,
$\left|\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\cosh\!\left(\displaystyle\frac{2n-1}{2}y\right)}\right|\leq\displaystyle\frac{(2n-1)^{p}}{\cosh\!\left(\displaystyle\frac{2n-1}{2}\delta\right)}\leq 2(2n-1)^{p}e^{-\tfrac{2n-1}{2}\delta},$
hence $\widetilde{G}_{p,0,1}(y)$ converges uniformly on $\delta,\infty)$ by the Weierstrass M-test. Moreover, differentiating
$f_{n}(y):=\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\cosh\!\left(\displaystyle\frac{2n-1}{2}y\right)}\quad\text{gives}\quad f_{n}^{\prime}(y)=-\displaystyle\frac{(-1)^{n}(2n-1)^{p+1}}{2}\,\displaystyle\frac{\sinh\!\left(\displaystyle\frac{2n-1}{2}y\right)}{\cosh^{2}\!\left(\displaystyle\frac{2n-1}{2}y\right)}.$
Using $\bigl|\sinh t\bigr|/\cosh^{2}t\leq 2e^{-t}$ for $t\geq 0$, we obtain for $y\geq\delta$,
$|f_{n}^{\prime}(y)|\leq C(2n-1)^{p+1}e^{-\tfrac{2n-1}{2}\delta},$
so $\sum_{n\geq 1}f_{n}^{\prime}(y)$ converges uniformly.
Differentiating $\tilde{G}_{p-2,0,1}\left(y\right)$ yields
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\tilde{G}_{p-2,0,1}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}{\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{(-1)^{n}\left(2n-1\right)^{p-2}}{\cosh\left(\displaystyle\frac{2n-1}{2}y\right)}=-\displaystyle\frac{1}{2}}\tilde{G}_{p-1,-1,2}\left(y\right).$
Using the identity $\frac{\mathrm{d}x}{\mathrm{d}y}=-x(1-x)z^{2}$ (see [5, P. 120, Entry. 9(i)] or (2.11)), we obtain
$\tilde{G}_{p-1,-1,2}\left(y\right)=2x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\tilde{G}_{p-2,0,1}\left(y\right).$ (5.70)
Applying Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we deduce
@@
Consequently, (5.69) and (5.70) can be rephrased as
$\tilde{G}_{p-2,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)=(-1)^{(p-1)/2}\left(\displaystyle\frac{yz}{\pi}\right)^{p+1}\displaystyle\frac{\mathrm{P}_{p-2}\left(1-x\right)}{2}\sqrt{x(1-x)}$ (5.71)
@@
$\tilde{G}_{p-1,-1,2}\left(\displaystyle\frac{\pi^{2}}{y}\right)$ $=2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}\left(1-x\right)}\tilde{G}_{p-2,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$=-2x\left(1-x\right)\left(\displaystyle\frac{yz}{\pi}\right)^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}\tilde{G}_{p-2,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.72)
@@
By evaluating specific instances of the right-hand side of (5.5) using Mathematica, we obtain
@@
$\Big.\quad+\left(-79360x^{5}+198400x^{4}-166112x^{3}+50768x^{2}-3698x+1\right)\sqrt{1-x}z\Big$.
$\Big.\quad+50521x^{5}-138933x^{4}+130250x^{3}-45530x^{2}+3693x-1\Big\},$
@@
$\Big.\quad+\left(-313155584x^{7}+1096044544x^{6}-1458129408x^{5}\atop+905212160x^{4}-255034240x^{3}+25361472x^{2}-298946x+1\right)\sqrt{(1-x)}z\Big$.
$\Big.\quad+(199360981x^{7}-747603679x^{6}+1074680289x^{5}-728130163x^{4}\atop+226132303x^{3}-24738669x^{2}+298939x-1)\Big\}.$
Setting p=4m1p=4m-1 and $x=\tfrac{1}{2}$ in Theorem 5.5, and noting that $\mathrm{P}^{\prime}_{4m-3}(1/2)=0$ (see the proof of Theorem 3.7 in [19]), we obtain the following corollary.
@@
Set $\Gamma=\Gamma(1/4)$. For any integer $m>0$ we know that
@@
$\quad+\displaystyle\frac{1}{2^{4m}}(4m-2)\mathrm{P}_{4m-3}\left(\displaystyle\frac{1}{2}\right)\displaystyle\frac{\Gamma^{8m-4}}{\pi^{6m-2}}.$ (5.73)
@@
Set $\Gamma=\Gamma(1/4)$. By letting $y=\pi$ and $x=1/2$ in Example $5.6$ with the help of Mathematica, we arrive at the example below
@@
Let $p\geq 3$ be an odd integer. We obtain
$\tilde{G}_{p,2,1}\left(y\right)$ $=\displaystyle\frac{1}{2}z^{p+1}\mathrm{P}_{p}(1-x)\sqrt{x(1-x)}-pz^{p+1}z^{\prime}x(1-x)\mathrm{A}_{p-1}(1-x)\sqrt{x}$ (5.74)
@@
where $\mathrm{P}_{p}\left(x\right)$ and $\mathrm{A}_{p-1}\left(x\right)$ are the coefficient polynomials from the Maclaurin expansions of $\mathrm{sd}\left(u\right)$ and $\mathrm{nd}\left(u\right)$ respectively.
@@
We initiate our derivation with equation (4.44), which gives
$\tilde{G}_{p,2,1}\left(y\right)$ $=-(-1)^{(p-1)/2}\displaystyle\frac{\pi^{p+1}}{y^{p+1}}\tilde{G}_{p,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)-(-1)^{(p-1)/2}\displaystyle\frac{p2^{p}\pi^{p}}{y^{p}}G_{p-1,0,1}\left(\displaystyle\frac{\pi^{2}}{y}\right)$
$\quad+(-1)^{(p-1)/2}\displaystyle\frac{2^{p}\pi^{p+2}}{y^{p+2}}G_{p,-1,2}\left(\displaystyle\frac{\pi^{2}}{y}\right).$ (5.75)
The derivative of $G_{p-1,0,1}\left(y\right)$ satisfies
$\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}G_{p-1,0,1}\left(y\right)=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{\mathrm{d}}{\mathrm{d}y}\displaystyle\frac{(-1)^{n}n^{p-1}}{\cosh(ny)}=-G_{p,-1,2}\left(y\right).$
From the fundamental relation $\frac{\mathrm{d}x}{\mathrm{d}y}=-x(1-x)z^{2}$ ([5, P. 120, Entry. 9(i)] or [19]), we have
$G_{p,-1,2}\left(y\right)=x\left(1-x\right)z^{2}\displaystyle\frac{\mathrm{d}}{\mathrm{d}x}G_{p-1,0,1}\left(y\right).$ (5.76)
Utilizing Lemma 3.2, for $\Omega\left(x,e^{-y},z,z^{\prime}\right)=0$, we obtain